OpenStudy (anonymous):

You have a pizza with a 16 inch diameter cut into 16 equal slices. A. What is the side length of a piece of pizza? B. What is the arc measure of a single slice?

OpenStudy (kc_kennylau):



OpenStudy (paki):

here Pi * r ^2 Pi=~3.14 r=radius=diameter/2=16/2= 8... now put in formula... 3.14*8^2=201.061 now divide this into the required number of pieces... which is 16... so 201.061/16 of the pizza= 12.56 inches

OpenStudy (paki):

and for arc measure of a single slice... we know that the pizza is circular having total circle length and total angles are 360... so just divide it by the nos... here it is... 360/16= 22.5 degrees...

OpenStudy (anonymous):

Thank you!

OpenStudy (paki):

most welcome...

OpenStudy (dwelsher):

c=2pi*r 2r=d ?? r=?